OpenStudy (anonymous):

@Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 @ganeshie8 @AravindG @robtobey @sauravshakya @thomaster @Nurali @hba @Yahoo! @Luigi0210 @dan815 @jdoe0001 @CliffSedge @kelliegirl33 @ivettef365 @Loser66 @SWAG @nubeer @DLS

OpenStudy (luigi0210):

Either use elimination or substitution, either works.

OpenStudy (anonymous):

(4, -6)

OpenStudy (anonymous):

I want to use elimination but I dont know it works

OpenStudy (texaschic101):

elimination... if you add the two equations, you will see that the y's cancel each other out x - y = 10 2x + y = 2 ----------------add 3x = 12 x = 4 now sub 4 in for x in either of the original equations x - y = 10 4 - y = 10 -y = 10 - 4 -y = 6 y = -6 check... 2x + y = 2 2(4) + (-6) = 2 8 - 6 = 2 2 = 2 (correct) x = 4 and y = -6 or (4,-6) any questions ?

OpenStudy (anonymous):

x - y = 10 x-10 = y 2x + y = 2 2x +(x-10) = 2 3x = 12 x = 4 x-10 = y 4-10 = y -6 = y